# Unlock the Power of Frozen LLMs in Knowledge Graph Completion

### v1

@@ 3.3 Main Results (For RQ1)
LLAMA-SVM 0.824 0.869 0.599
LLAMA-LR 0.840 0.863 0.665
LLAMA-MLP 0.851 0.874 0.679
@@
LLAMA-MLP 0.851 0.874
LLAMA-DES-H 0.895 -
LLAMA-DES2 (GPT)* 0.890 0.892
@@
MISTRAL-MLP 0.848 0.903
MISTRAL-DES-H 0.884 -
MISTRAL-DES2 (GPT)* 0.875 0.912
@@
GEMMA-MLP 0.810 0.895
GEMMA-DES-H 0.895 -
GEMMA-DES2 (GPT)* 0.851 0.911
@@
3.4 Ablation Study (For RQ2)
Through the analysis depicted in the left portion of Figure 3, it is observed that with an increase in the index of the intermediate layer, the efficacy of hidden states for triple classification demonstrates an initial incline followed by a subsequent decline. Intermediate layers closer to natural language text output generally exhibit better prediction performance compared to the lowest layers. This observation corroborates the notion posited by Geva et al. (2020) that language models encode knowledge differentially across their Feed-Forward layers, with higher layers assimilating information from lower layers. During pretraining, it is possible that lower layers may not have stored certain knowledge, hence they cannot generate hidden states with effective triple classification effects. Furthermore, the prediction performance of the layers closer to the natural language text output is generally worse than those in the middle layers. This phenomenon may be attributed to the hallucination issue inherent in LLaMA, wherein higher layers could potentially either omit or alter triple information, leading to a blurred distinction between positive and negative triple hidden states induced by stimuli. However, the lower layers of LLaMA can effectively gather the stimulus information required for the completion of the knowledge graph tasks. Akin observations are discernible on the right side of Figure 3 for the Mistral model, which has a different backbone from LLaMA. Meanwhile, as observed in Table 5, our method demonstrates strong versatility across different backbone models, achieving performance comparable to full dataset fine-tuning with only a small sample size.
Regarding classification models, a comparative analysis encompassing MLP, SVM, and Logistic Regression, as delineated in Table 4, is conducted. The result shows that MLP outperforms all three datasets compared to the other two classification models, while Logistic Regression also exhibits results close to MLP. This underscores the efficacy of our stimulus methodology, whereby hidden states of positive and negative samples within a high-dimensional space exhibit linear separability.
@@ 3.5 Efficiency Study (For RQ3)
From Figure 4, we can observe the following: For the FB13 dataset, using 400 training samples, which is 0.06% of the training set, achieves 98.3% of the performance of using the entire training set; For the FB15K-237N dataset, using 0.57% of the training set achieves 99.6% of the performance of using the entire training set; For the WN18RR dataset, using 0.46% of the training set achieves 98.8% of the performance of using the entire training set; For the UMLS dataset, using 9.58% of the dataset achieves 88.7% of the performance of using the entire training set. Our experimental results confirm that the language model is capable of generating effective hidden states through the techniques described in Sections 2.2 and 2.3, demonstrating its potential for data-efficient scenarios within the context of knowledge graph completion tasks.
Meanwhile, we also evaluate the compute efficiency of our method. Table 8 shows that our method can achieve significantly lower consumption compared to fine-tuning even in the setting of the full dataset. Compared to fine-tuned LLMs, we boost the GPU memory efficiency to 188×188$\times$188 × during model training and speed up 13.48×13.48$\times$13.48 × overall. Please refer to Appendix B.3 for details.
@@ 4.1 Knowledge Graph Completion
Knowledge graphs are typically characterized by their inherent incompleteness, often containing substantial amounts of implicit or missing valuable knowledge (Hogan et al., 2021). The prevailing methodologies in KGC can be categorized into two main classes: structure information-based methods and additional information-based methods (Shen et al., 2022b). Structure information-based methods leverage the inherent structural data present within knowledge graphs. For example, RESCAL (Nickel et al., 2011), DistMult (Yang et al., 2014), COMPGCN (Vashishth et al., 2019), and ConvE (Dettmers et al., 2018) primarily use scoring functions to evaluate the semantic similarities of entity or relation embeddings within a latent embedding space to infer missing information. Additional information-based methods incorporate a variety of supplementary data to enrich the knowledge graphs. These methods utilize node attributes, entity-related information, or relational path data to enhance the features of knowledge embeddings critical for KGC tasks. The prominent examples are KG-BERT (Yao et al., 2019) and LASS (Shen et al., 2022a), which assess the plausibility of triples by transforming the triple and its contextual data into natural language sentences using predefined templates. These sentences are then encoded using a fine-tuned BERT model, effectively leveraging additional information to improve the accuracy of triple classification. KG-S2S (Chen et al., 2022) and KGT5 (Saxena et al., 2022) employ encoder-decoder architecture language models to complete KGC in a generative manner. KG-LLAMA (Yao et al., 2023) utilizes triples composed of entity and relation names as sentences, fine-tuning the LLaMA model for KGC using natural language. Li et al. (2024) enhances KGC with LLM by generating entity descriptions and designing loss functions for fine-tuning.
@@ 5 Conclusion
In this paper, we introduce a novel data-efficient probing method that completes KGC by leveraging the internal knowledge of LLMs. Experiments show that using the probing method with generated entity descriptions, frozen LLMs can achieve performance comparable to that of their fine-tuned counterparts. Simultaneously, this method saves a significant amount of GPU memory and time compared to fine-tuning approaches. This work explores the potential of frozen LLMs in KGC tasks with both efficacy and efficiency. Compared to fine-tuning LLMs, we hit the core of KGC problem in a better way.
@@
In this study, our method was primarily tested using the UMLS domain-specific knowledge graph and was not applied to other domain-specific KGs. Our current architectural design is specifically tailored for triple classification and relation prediction within knowledge graph completion. Recognizing these limitations, future work will aim to expand our architecture to accommodate additional KGC tasks such as link prediction. This extension will potentially broaden the applicability and enhance the robustness of our method across a more diverse set of knowledge graph scenarios.
@@ A.2 Implementation Details
In our implementation, we use the grid search to find appropriate sets of hyperparameters for baselines and our proposed method.
For the MLP classifier configuration, we opt for a batch size of 64, a learning rate of 3e-5, and utilize the AdamW optimizer, with training restricted to 30 epochs. For LLaMA-7B-ICL and Alpaca-7B-ICL, we randomly select one positive sample and one negative sample from the training set of each dataset, respectively. The prompt for KGC using Alpaca-7B-ICL on the WN18RR dataset is illustrated in Figure 6. Additionally, the hyper-parameters for models trained with the LoPA technique, specifically KG-LLAMA-7B, are outlined in Table 7.
@@ B.3 Efficiency Analysis
B.4 Analysis of Methods for Entity Description Generation
@@
B.5 Analysis of Using Frozen LLM in KGC
It has been observed that the predictive performance achieved by the frozen language models can not achieve comparable performance with supervised fine-tuning the language models (Zhang et al., 2023; Yao et al., 2023). But the knowledge graphs used in the experiments are encyclopedic knowledge graphs, such as FB15K-237 (Toutanova and Chen, 2015), which are subsets of FB15k (Bordes et al., 2013) derived from Wikipedia. Large Language Models like LLaMA (Touvron et al., 2023), which are pretrained on datasets such as Wikipedia, CommonCrawl, etc., have stored this type of world knowledge (Davison et al., 2019). Liu et al. (2023) shows that prompting can effectively access the knowledge within LLMs without fine-tuning. Therefore, frozen LLMs have sufficient knowledge and capability to complete KGC tasks.

### v2

@@ 3.4 Ablation Study (For RQ2)
LLAMA-SVM 0.821 0.866 0.594
LLAMA-LR 0.839 0.860 0.662
LLAMA-MLP 0.855 0.872 0.674
@@
LLAMA-MLP 0.855 0.872
LLAMA-DES-H 0.901 -
LLAMA-DES2 (GPT)* 0.912 0.917
@@
MISTRAL-MLP 0.846 0.901
MISTRAL-DES-H 0.880 -
MISTRAL-DES2 (GPT)* 0.893 0.925
@@
GEMMA-MLP 0.806 0.893
GEMMA-DES-H 0.891 -
GEMMA-DES2 (GPT)* 0.867 0.917
@@ 3.5 Efficiency Study (For RQ3)
As illustrated in Figure 4, our method notably surpasses both KG-LLAMA and KG-BERT on the FB13 when trained with only 100 samples, as opposed to the 3000 samples used by the latter methods. For the UMLS, the performance of KG-LLAMA and KG-BERT with 3000 training samples is comparable to that of LLAMA-DES-H trained with only 500 samples. Additionally, combining the results from Table 2 and Table 10, for the FB13 dataset, our method attains 97.2% of the performance achievable with the full training set, utilizing only 500 samples, which constitutes a mere 0.079% of the total training samples. In contrast, KG-LLAMA and KG-BERT achieve only 78.7% and 63.1% of the maximum performance, respectively. For the UMLS, with 9.58% of the training set, our method reaches 90.6% of the full training set performance. For other datasets, please refer to Appendix B.6. Our experimental results confirm that the language model is capable of generating effective hidden states through the techniques described in Sections 2.2 and 2.3, demonstrating its potential for data-efficient scenarios within the context of KGC tasks.
Figure 4: The prediction performance of KG-LLAMA, KG-BERT, and LLAMA-DES-H on triple classification varies with the size of the training dataset. The horizontal axis represents the number of training samples used, with an equal number of positive and negative samples.
Meanwhile, we also evaluate the compute efficiency of our method. Table 8 shows that our method can achieve significantly lower consumption compared to parameter-efficient fine-tuning even in the setting of the full dataset. Compared to fine-tuned LLMs, we boost the GPU memory efficiency to 188×188$\times$188 × during model training and speed up 13.48×13.48$\times$13.48 × overall. Please refer to Appendix B.3 for details.
@@ 4.1 Knowledge Graph Completion
Knowledge graphs are typically characterized by their inherent incompleteness, often containing substantial amounts of implicit or missing valuable knowledge (Hogan et al., 2021). The prevailing methodologies in KGC can be categorized into two main classes: structure information-based methods and additional information-based methods (Shen et al., 2022b). Structure information-based methods leverage the inherent structural data present within knowledge graphs. For example, RESCAL (Nickel et al., 2011), DistMult (Yang et al., 2014), COMPGCN (Vashishth et al., 2019), and ConvE (Dettmers et al., 2018) primarily use scoring functions to evaluate the semantic similarities of entity or relation embeddings within a latent embedding space to infer missing information. Additional information-based methods incorporate a variety of supplementary data to enrich the knowledge graphs. These methods utilize node attributes, entity-related information, or relational path data to enhance the features of knowledge embeddings critical for KGC tasks. The prominent examples are KG-BERT (Yao et al., 2019) and LASS (Shen et al., 2022a), which represent entities, relations, and triples as textual sequences, turning KGC into a sequence classification problem. KG-S2S (Chen et al., 2022) and KGT5 (Saxena et al., 2022) employ encoder-decoder architecture language models to complete KGC in a generative manner. KG-LLAMA (Yao et al., 2023) utilizes triples composed of entity and relation names as sentences, parameter-efficient fine-tuning the LLaMA model for KGC. Li et al. (2024) enhances KGC with LLM by generating entity descriptions and designing loss functions for fine-tuning.
@@ 5 Conclusion
In this paper, we introduce a novel data-efficient probing method that completes KGC by leveraging the internal knowledge of LLMs. Experiments show that using the probing method with generated entity descriptions, frozen LLMs can achieve performance comparable to that of their fine-tuned counterparts. Simultaneously, this method saves a significant amount of GPU memory and time compared to parameter-efficient fine-tuning approaches. This work explores the potential of frozen LLMs in KGC tasks with both efficacy and efficiency. Compared to fine-tuning LLMs, we hit the core of KGC problem in a better way.
@@
Our current architectural design is only applicable to open-source LLMs or closed-source models that provide access to token embeddings; it is not suitable for closed-source models that exclusively output text. Recognizing these limitations, future work will aim to expand our architecture to accommodate closed-source LLMs that only output text or token probabilities. Additionally, future efforts could integrate structure information-based KG models, such as TransE, with our methodology in a hybrid approach.
@@ A.2 Implementation Details
In our implementation, we experimented with these baselines using the hyperparameters recommended in their paper.
For the negative sampling employed on the datasets, we utilize the same negative sampling technique as Yao et al. (2019) for FB13 and WN11. Specifically, FB15K-237N is designed to extract hard negative triples, providing a more challenging evaluation and reducing false negatives in the validation/test datasets during dataset construction. A similar approach is adopted for constructing negative samples for WN18RR and UMLS. Meanwhile, for all experiments in our study, the training sets contain an equal number of positive and negative samples.
For the MLP classifier configuration, we opt for a batch size of 64, a learning rate of 3e-5, and utilize the AdamW optimizer, with training restricted to 30 epochs. For LLaMA-7B-ICL and Alpaca-7B-ICL, we randomly select one positive sample and one negative sample from the training set of each dataset, respectively. The prompt for KGC using Alpaca-7B-ICL on the WN18RR dataset is illustrated in Figure 6. Additionally, the hyper-parameters for models trained with the LoRA technique, specifically KG-LLAMA-7B, are outlined in Table 7.
@@ B.3 Efficiency Analysis
Method Pred Samples FB13 WN11 15K237N WN18RR UMLS
LLAMA-7B N-L 0 0.699 0.661 0.573 0.458 0.658
ALPACA-7B N-L 0 0.226 0.696 0.561 0.714 0.526
LLAMA-7B-ICL N-L 2 0.501 0.500 0.578 0.502 0.538
ALPACA-7B-ICL N-L 2 0.675 0.706 0.607 0.691 0.605
STRUCTURAL-AWARE-IT N-L ALL 0.907 0.959 0.752 0.927 0.881
ALPACA-DES N-L 0 0.611 - 0.590 0.548 0.525
LLAMA-MLP H-MLP 3k/10k 0.855 0.872 0.674 0.858 0.882
LLAMA-DES-H H-MLP 3k/10k 0.901 - 0.738 0.931 0.862
LLAMA-DES2 (GPT)* H-MLP 3k/10k 0.912 0.917 0.726 0.924 0.860
LLAMA-DES-H H-MLP ALL 0.905 - 0.742 0.937 0.866
Table 10: The experiment results (accuracy) of triple classification. ALPACA-DES combines non-generated descriptions with $PT_{2}$ for Alpaca-7B.
B.4 Additional Baselines
From Table 10, it is noteworthy that ALPACA-DES fails to outperform ALPACA-7B, suggesting that mere incorporation of entity descriptions may not yield direct performance enhancements, considering the possibility of LLMs revealing the wrong responses and being influenced by hallucinations.
Meanwhile, in our study, we did not compare our method with KOPA because it integrates KG models (such as TransE/RotatE) with language models (like LLaMA) in a hybrid approach, thereby blending multiple approaches rather than solely utilizing the capabilities of LLMs (Zhang et al., 2023). In contrast, our current research exclusively focuses on the utilization of language models (like LLaMA), rendering KOPA an unsuitable baseline for our purposes. Nonetheless, we reproduced the Structural-aware-IT model proposed by Zhang et al. (2023), which incorporates in-context learning and instruction tuning. This model samples the neighborhoods of the head entity $h$ and tail entity $t$, including the textual descriptions of neighboring triples in the demonstration prompt, followed by fine-tuning on the entire training set. Like KOPA, this approach integrates structural information from the KG with LLM capabilities. Despite this, our method, which exclusively utilizes LLMs without integrating KG structural information, achieves performance comparable to the Structural-aware-IT model, which combines both KG structure and LLM capabilities.
B.5 Analysis of the Performance in the Intermediate Layers of LLM
The selection of the optimal layer index in LLMs serves as a hyperparameter, which we may select based on empirical evidence or adjust as needed when a validation set is available. Meanwhile, our experiments found a strong consistency for the same LLM across different knowledge graphs (for example, for LLAMA, the 15th layer often yields the best KGC performance), as shown in Figure 3, which reduces the complexity of utilization.
Additionally, the superior performance of the intermediate layers of language models, as compared to those closer to the output layer, may be attributed to the potential for each layer of the model to experience hallucinations. Moreover, each layer of a language model can potentially propagate hallucinations from previous layers, manifesting as deliberate falsehoods or the expression of misleading information (Zou et al., 2023). This phenomenon can lead to scenarios where models often know the correct answer even while generating incorrect outputs (Zou et al., 2023). Concurrently, our experimental results are consistent with the conclusions of Jiang et al. (2024), which found that the MLP layers in the later stages have an inhibitory effect on the extracted knowledge, particularly contributing significantly to erroneous output decoding. As a result, the middle layers in their experiments also tend to outperform the later layers.
B.6 Additional Analysis of the Data Efficiency
From Figure 8, we can observe the following: For the FB13 dataset, using 400 training samples, which is 0.06% of the training set, achieves 97.0% of the performance of using the entire training set; For the FB15K-237N dataset, using 0.57% of the training set achieves 99.6% of the performance of using the entire training set; For the WN18RR dataset, using 0.46% of the training set achieves 98.8% of the performance of using the entire training set; For the UMLS dataset, using 9.58% of the dataset achieves 90.6% of the performance of using the entire training set.
It can be observed that the LLAMA-DES-H method exhibits slightly lower data efficiency on the UMLS dataset compared to other methods. This phenomenon may be attributed to the fact that UMLS is a domain-specific knowledge graph, whereas the other datasets are encyclopedic knowledge graphs. The increased complexity of the decision boundary in the hidden states of the positive and negative example triplets generated by LLaMA necessitates a larger volume of data to adequately model the boundary.
B.7 Analysis of Methods for Entity Description Generation
@@
B.8 Analysis of Using Frozen LLM in KGC
It has been observed that the predictive performance achieved by the frozen language models can not achieve comparable performance with supervised fine-tuning the language models (Zhang et al., 2023; Yao et al., 2023). But the knowledge graphs used in the experiments are encyclopedic knowledge graphs, such as FB15K-237 (Toutanova and Chen, 2015), which are subsets of FB15k (Bordes et al., 2013) derived from Wikipedia. LLMs like LLaMA (Touvron et al., 2023), which are pretrained on datasets such as Wikipedia, CommonCrawl, etc., have stored this type of world knowledge (Davison et al., 2019). Liu et al. (2023) shows that prompting can effectively access the knowledge within LLMs without fine-tuning. Therefore, frozen LLMs have sufficient knowledge and capability to complete KGC tasks.